OpenStudy (anonymous):

how to find determinant of Matrices A = {{1,-2,3}{2,1,-1}{1,0,2}} with Elementary Row Operations methode

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

ganeshie8 (ganeshie8):

\[ \left| \begin{array}{ccc} 1 & -2 & 3 \\ 2 & 1 & -1 \\ 1 & 0 & 2 \end{array} \right| \]

ganeshie8 (ganeshie8):

since 0 is there in 32, lets try to expand the determinant wid that column, before that, lets make another element in that column 0

OpenStudy (anonymous):

http://www.wikihow.com/Find-the-Determinant-of-a-3X3-Matrix If u want to learn to do it in only 6 steps

ganeshie8 (ganeshie8):

R1+2R2 \[ \left| \begin{array}{ccc} 5 & 0 & 1 \\ 2 & 1 & -1 \\ 1 & 0 & 2 \end{array} \right| = 10 - 1 = 9 \]

OpenStudy (anonymous):

i mean the Elementary Row Operations method :)

ganeshie8 (ganeshie8):

yup ! we did one elementary row operation R1+2R2

ganeshie8 (ganeshie8):

you want to convert the whole matrix to upper triangular matrix, and multiple the diagonal is it ?

OpenStudy (anonymous):

yaph that's what i mean :)

ganeshie8 (ganeshie8):

cool, then convert it to upper triangular matrix ? :)

ganeshie8 (ganeshie8):

\( \left| \begin{array}{ccc} 1 & -2 & 3 \\ 2 & 1 & -1 \\ 1 & 0 & 2 \end{array} \right| \)

ganeshie8 (ganeshie8):

R2-2R1 R3-R1 \( \left| \begin{array}{ccc} 1 & -2 & 3 \\ 0 & 5 & -7 \\ 0 & 2 & -1 \end{array} \right| \)

ganeshie8 (ganeshie8):

wat next ?

OpenStudy (anonymous):

i don't understand hoe to make it into upper triangular matrice -_-

ganeshie8 (ganeshie8):

so far, u okays ? :)

OpenStudy (anonymous):

okay so far i understand )

OpenStudy (anonymous):

what next ?

ganeshie8 (ganeshie8):

good :), upper triangular means, the matrix must have 0s in the lower triangle region like below \( \left| \begin{array}{ccc} a & b & c \\ 0 & d & e \\ 0 & 0 & f \end{array} \right| \)

ganeshie8 (ganeshie8):

our matrix still missing one 0, right ? at 32 position we have 2, lets make it 0

OpenStudy (anonymous):

yaph there is still 1 missing zero okay so far so good :)

ganeshie8 (ganeshie8):

\( \left| \begin{array}{ccc} 1 & -2 & 3 \\ 0 & 5 & -7 \\ 0 & 2 & -1 \end{array} \right| \)

ganeshie8 (ganeshie8):

how to make it 0 ? will below work ? R3 - (2/5)R2

ganeshie8 (ganeshie8):

R3 - (2/5)R2 \( \left| \begin{array}{ccc} 1 & -2 & 3 \\ 0 & 5 & -7 \\ 0 & 0 & 9/5 \end{array} \right| \)

ganeshie8 (ganeshie8):

now its complete! to find det, just multiply the diagonal elements

OpenStudy (anonymous):

wow you're so cool :D thank you very much i really confused how to solve this before, and another question when we make the equation like R1+2R2 are we make it by trying and error or there is special terms ?

ganeshie8 (ganeshie8):

np :) no trial and error, matlab does this in a systematic way

ganeshie8 (ganeshie8):

\( \left| \begin{array}{ccc} 1 & -2 & 3 \\ 2 & 1 & -1 \\ 1 & 0 & 2 \end{array} \right| \)

ganeshie8 (ganeshie8):

thats the given matrix, First i want to make 21 element 0, i see that 2 is sitting there. so, to make it 0 using row operations, i have to R2-2R1

ganeshie8 (ganeshie8):

Next i want to make 31 element also 0, i see that 1 is sitting there. so, to make it 0 using row operations, i have to do R2-R1

ganeshie8 (ganeshie8):

Notice that, I am subtracting multiples of R1 in both cases; 11 element is our pivot when making 21 and 31 elements 0

OpenStudy (anonymous):

i see so we find the equation by seeing the 21, 31 and 32 right ? okay thank you :)

ganeshie8 (ganeshie8):

exactly !, and in that order